A car of mass 500kg accelerates from rest at 1m/s². What is the total distance covered in 1 minute?
1
Expert's answer
2019-04-04T09:06:00-0400

We can find the total distance covered in 1 minute from the kinematic equation:


"s = v_0t + \dfrac{1}{2}at^2,"

here, "s" is the total distance covered in 1 minute, "v_0 = 0" is the initial velocity of the car, "a" is the acceleration of the car and "t" is time.

Then, we get:


"s = v_0t + \dfrac{1}{2}at^2 = 0 \dfrac{m}{s} \cdot 60 s + \dfrac{1}{2} \cdot 1 \dfrac{m}{s^2} \cdot (60 s)^2 = 1800 m."

Answer:

"s = 1800 m."